# Facial Surgery Preview Based on the Orthognathic Treatment Prediction

### v1

@@ 4.1 Data Collection and Pre-processing
Automatic Landmarks annotation. We used Supervision by Registration [9] with a facial landmark detector to annotate landmarks. A point light was placed in front of the patient’s face to render the mesh to a frontal view image. We then applied a pre-trained landmark detector to extract $2$D coordinates for $68$ landmarks and transformed them back to 3D using the vertex-to-pixel record. Any drift in landmarks for challenging preoperative cases was manually corrected.
4.2 Training details
@@
The training process took approximately 25 minutes to complete. To ensure the robustness and reliability of our results, we employed a 5-fold cross-validation strategy. The 163 valid pairs of facial scans were shuffled and split into 5 consecutive folds. Each fold was used once as the validation set, while the remaining four folds formed the training set. Data splitting was performed prior to data augmentation, which was applied exclusively to the training dataset. We report the average score in Table 2.
4.3 Gallery
@@
In our gallery, we present four patients with their real postoperative appearances alongside prediction results from our algorithm and the linear network using the Basel model. To assess accuracy, we also conducted texture transfer experiments on the Basel network, comparing it to LARS, the best model from [12]. Bimaxillary orthognathic surgery, with or without genioplasty, was performed for all four patients, shown in Fig. 3. For the first patient, our approach effectively predicted the outcome of mouth correction, with minor discrepancies in the lower face compared to LARS. In the lower nasal region, the linear model predicted a longer, wider, and lower nasal base, deviating from the surgical plan. For the second patient, our algorithm successfully predicted the jaw’s repositioning, while LARS did not resolve the protruding chin. For the third and fourth patients, both of whom had bimaxillary orthognathic surgery with genioplasty to shorten facial length by adjusting the chin’s tilt angle, our algorithm accurately predicted the jaw’s angle change, resulting in a more aesthetically pleasing facial shape. LARS, however, failed to predict these changes accurately.
4.4 ablation studies
@@
We conducted an ablation study to examine the effectiveness of the four losses and data augmentation techniques introduced in our model. Both quantitative (as shown in Table. 1) and qualitative (as shown in Figure. 4) results demonstrate the efficacy of these techniques. Our findings suggest that the incorporation of these four losses and data augmentation can significantly improve the performance and efficacy of our model.
@@
Figure 3: Comparison of our method with LARS across four patient cases. Ground truth (GT) pre- and post-surgical scans are shown on the left for reference. LARS predictions overestimated nose size and face length while underestimating lip thickness in the first, third, and fourth cases, and failed to adequately address the prominent chin in the second case. Our method effectively predicts surgical outcomes while preserving the patient’s natural appearance.
Figure 4: Removing the mouth-convexity loss leads to an overly prominent chin, while removal of the asymmetry loss results in facial asymmetry. Without the latent code loss, there is significant drift in the cheeks, and the overall error increases when the geometry loss is excluded. The addition of data augmentation can further reduces the prediction error.
4.5 User study
The user experiment aimed to test whether participants could distinguish between machine learning-generated faces and real surgical outcomes. We collected pre- and post-surgery 3D data for each patient, parameterized with FLAME and transferred textures for consistency. To aid participants in observing facial symmetry and feature positioning, we provided frontal and side views, along with an animation rotating 180 degrees from left to right. Preoperative data was clearly labeled, and participants had to identify if the real surgical outcome appeared before or after the machine-predicted result. Participants included five medical professionals (two surgeons, three orthodontists) and 15 engineers ($12$ males, $3$ females). Using a $2$ Alternative Forced Choice format and a t-test, results showed $49.3\%$ of medical professionals and $50.2\%$ of engineers chose our solution, with p-values of $0.84$ and $0.88$, respectively. This indicated neither group could distinguish between the two outcomes.
4.6 Quantitative comparison
We quantitatively compared the error of our prediction network with that of LARS, and found that our network outperformed LARS in both the Chamfer distance and Hausdorff distance metrics, as shown in Table 2. Additionally, we observed that data augmentation was very helpful in improving our performance, but had little effect on the performance of LARS.
Table 2: Performance Comparison
Algorithms HD* (mm) $\downarrow$ CD* (mm) $\downarrow$ Data Amount
mean min max mean min max
OUR’s 9.00 7.63 11.30 2.50 1.24 3.60 1330
LARS 9.68 7.50 15.41 2.77 1.72 4.77 1330
OUR’s w/o synthesized data 9.43 8.00 13.53 2.94 1.91 6.38 133
LARS w/o synthesized data 9.67 7.46 16.13 2.98 1.57 5.21 133
  • 1.
    * HD and CD represent Hausdorff and Chamfer Distance respectively.
4.7 application
@@ 4.6 Quantitative comparison
We created an animation by interpolating the latent vectors, which can be used to show patients a transformation from pre-surgery to post-surgery, as shown in Fig. 5. This allows patients to visualize the movement of their chins and the change in their overall facial appearance. For a better demonstration, please refer to our supplementary video.
5 Conclusion
In this paper, we introduce a novel method for predicting facial appearance following orthognathic treatment using only 3D face geometry. During the training phase, our approach utilizes customized mouth-convexity and asymmetry losses, combined with latent code and geometric losses, to outperform existing methods in terms of accuracy. Our experimental results demonstrate significant potential for real-world clinical applications. Validating the method’s effectiveness in actual clinical settings is planned for our future research.
Limitation. Our current model doesn’t offer multimodal predictions based on factors like age, gender, and skin condition due to the use of unlabeled data. However, with additional multi-modal labeled data in the future, we can categorize the existing data and make more targeted predictions without significantly reducing the dataset. Our focus is on providing automatic visualizations for patients, not orthognathic professionals. In the future, we may add user-friendly interfaces with adjustable parameters for medical professionals if we develop a platform for surgical planning.

### v2

@@ 3.2.2 Landmarks annotation
We utilized Supervision by Registration [28] in conjunction with a facial landmark detector to automatically annotate facial landmarks. A point light was positioned in front of the patient’s face to render the 3D mesh into a frontal view image. A pre-trained landmark detector was then applied to extract 2D coordinates for 68 facial landmarks. These 2D coordinates were subsequently transformed back to 3D using the vertex-to-pixel mapping. For challenging preoperative cases, any discrepancies or drift in the landmark positions were manually corrected by a professional to ensure the dataset’s accuracy and reliability.
3.2.3 Training settings
@@
The training process took approximately 25 minutes to complete. To ensure the robustness and reliability of our results, we employed a 5-fold cross-validation strategy. The 163 valid pairs of facial scans were subjected to random shuffle and then split into 5 consecutive folds for cross-validation. Each fold was used once as the validation set, while the remaining four folds formed the training set. Data splitting was performed prior to data augmentation, which was applied exclusively to the training dataset. We report the average score in Table 4.
3.3 Analysis of user study results
Table 1: Comparison of Sensitivity, Specificity, and Precision values for Engineers and Medical Professionals in distinguishing Machine Learning-Generated faces (MLG) and Real surgical outcomes (RSO)
Medical Professionals Engineers
Sensitivity 53.30% 54.20%
Specificity 45.30% 46.20%
Precision 49.40% 50.20%
The sensitivity, specificity, and precision values are presented in Table 1. The results indicated that engineers had a slightly higher sensitivity percentage (54.20%) compared to medical professionals (53.30%), suggesting that engineers were slightly better at identifying Machine Learning-Generated faces (MLG) and Real Surgical Outcomes (RSO) accurately. However, both groups exhibited low specificity percentages, with engineers at 46.20% and medical professionals at 45.30%, indicating challenges in distinguishing between MLG faces and RSO. In terms of precision, engineers had a slightly higher percentage (50.20%) compared to medical professionals (49.40%), suggesting a slightly higher accuracy for engineers in identifying MLG faces. Nevertheless, there was no statistically significant difference in the ability to differentiate between MLG faces and RSO within each group, with p-values of 0.567 for engineers and 0.256 for medical professionals (Table 2). Additionally, Table 3 presents the confusion matrix, illustrating the classification performance for both groups. Noteworthily, although the subjective measure employed in this study holds value in assessing participants’ ability to differentiate between MLG faces and RSO, it is important to acknowledge that subjective evaluation alone may not entirely validate clinical accuracy.
Table 2: Comparison of Engineers’ and Medical Professionals’ ability to distinguish between Machine Learning-Generated faces and Real surgical outcomes
Medical Professionals ($N_{med}=5$) $n=150$ Engineers ($N_{eng}=15$) $n=450$
MLG 34 (23%) 104 (23%)
MLG identified as RSO 41 (27%) 121 (27%)
RSO identified as MLG 35 (23%) 103 (23%)
RSO 40 (27%) 122 (27%)
95% CI 12.58 to 17.02 14.01 to 16.12
P-value 0.567 0.256
  • $N_{med}$, Number of medical professionals; $N_{eng}$, Number of engineers; n, number of responses;
  • MLG, Machine learning generated faces; RSO, Real surgical outcomes; CI, Confidence Interval;
  • McNemar chi-square test was performed unless otherwise mentioned.
    $p<0.05$ (in bold italics), considered statistically significant.
Table 3: Confusion matrix presenting True Positives, False Positives, True Negatives, and False Negatives.
True Responses
RSO MLG
Predicted Responses Predicted RSO_med 53.30% 54.70%
Predicted MLG_med 46.70% 45.30%
Predicted RSO_eng 54.20% 53.80%
Predicted MLG_eng 45.80% 46.20%
  • RSO: Real surgical outcomes;
  • MLG: Machine learning-generated faces.
3.4 Quantitative metrics
We introduce two key metrics to quantitatively evaluate the accuracy of the predictions: the Hausdorff distance and the Chamfer distance. These metrics are commonly used to assess the similarity between two point clouds or surfaces, providing insights into the geometric alignment of the predicted and ground truth meshes.
The Hausdorff distance measures the greatest of all the distances from a point in one set to the closest point in the other set [29]. It is defined as:
$d_{H}(A,B)=\max\left(\sup_{a\in A}\inf_{b\in B}\|a-b\|,\sup_{b\in B}\inf_{a\in A}\|b-a\|\right)$ (7)
where $A$ and $B$ are two point sets, and $\|a-b\|$ denotes the Euclidean distance between points $a$ and $b$. The Hausdorff distance is sensitive to outliers and emphasizes the maximum deviation between the two sets.
The Chamfer distance provides a more balanced measure of the overall geometric difference between two point clouds [30]. It is computed as:
$d_{C}(A,B)=\frac{1}{|A|}\sum_{a\in A}\min_{b\in B}\|a-b\|^{2}+\frac{1}{|B|}\sum_{b\in B}\min_{a\in A}\|b-a\|^{2}$ (8)
where $|A|$ and $|B|$ are the number of points in sets $A$ and $B$, respectively, and $\|a-b\|$ is the Euclidean distance between points $a$ and $b$. The Chamfer distance measures the average distance between the points in both sets, offering a less sensitive alternative to the Hausdorff distance for assessing surface matching.
Both of these metrics allow for a precise comparison of the predicted and actual 3D facial geometries, providing valuable insights into the performance of our model.
Figure 4: Comparison of our method with LARS across four patient cases. On the left, the ground truth (GT) pre- and post-surgical scans are shown for reference. The middle columns display our predicted results, and the right columns show the LARS model’s predictions. To compare the prediction errors with the real outcomes, heatmaps are provided showing the error distribution across facial regions, with the error bars located in the bottom-right corner.
3.5 Quantitative comparison
These examples compare the real postoperative appearances of patients with the prediction results from our algorithm and LARS, the state-of-the-art model from [4]. We quantitatively compared the prediction errors of our network with those of LARS using the Chamfer distance (CD) and Hausdorff distance (HD) metrics. Table 4 summarizes the results across different settings.
[…]
Table 5 presents the results of the statistical analysis using a t-test to compare our model with the LARS model based on two distance metrics: Hausdorff and Chamfer distances. The t-statistics and corresponding p-values indicate that both metrics show statistically significant differences between the models. Specifically, for the Hausdorff distance, our model demonstrated a t-statistic of 2.113 with a p-value of 0.039, which is below the 0.05 threshold for significance. Similarly, for the Chamfer distance, the t-statistic was 2.143 with a p-value of 0.036, also demonstrating a significant difference. These results suggest that our model outperforms the LARS model in terms of both distance metrics.
Table 4: Performance Comparison
Algorithms HD* (mm) $\downarrow$ CD* (mm) $\downarrow$ Data Amount
mean min max mean min max
OUR’s 9.00 7.63 11.30 2.50 1.24 3.60 1330
LARS 9.68 7.50 15.41 2.77 1.72 4.77 1330
OUR’s w/o synthesized data 9.43 8.00 13.53 2.94 1.91 6.38 133
LARS w/o synthesized data 9.67 7.46 16.13 2.98 1.57 5.21 133
    HD and CD represent Hausdorff and Chamfer Distance respectively.
Table 5: Statistical analysis of Hausdorff and Chamfer distances.
Comparison Metric t-statistic p-value Significance
Our model vs LARS model Hausdorff 2.113 0.039 Significant
Our model vs LARS model Chamfer 2.143 0.036 Significant
3.6 Qualitative Comparison
To qualitatively evaluate our algorithm, we selected representative cases for visualization, as shown in Fig. 4. These examples compare the real postoperative appearances of patients with the prediction results from our algorithm and LARS  [4]. To ensure a fair assessment of accuracy, experiments were also conducted using the Basel network and compared against LARS. The showcased patients underwent bimaxillary orthognathic surgery, with or without genioplasty.
@@ 3.6 Qualitative Comparison
For the first patient, our approach effectively predicted the correction of mouth alignment, showing only minor discrepancies in the lower face when compared to LARS. Notably, in the lower nasal region, the linear model predicted a longer, wider, and lower nasal base, deviating significantly from the surgical plan. For the second patient, our algorithm successfully predicted the repositioning of the jaw, whereas LARS struggled to address the protruding chin accurately. In the cases of the third and fourth patients, both of whom underwent bimaxillary orthognathic surgery combined with genioplasty to shorten facial length by adjusting the chin’s tilt angle, our algorithm precisely captured the jaw’s angle changes, resulting in a more aesthetically balanced facial shape. In contrast, LARS failed to accurately predict these adjustments, underscoring the superior predictive capability of our method.
4 Discussion
4.1 Ablation experiment
@@
We conducted an ablation study to examine the effectiveness of the four losses and data augmentation techniques introduced in our model. Both quantitative (as shown in Table. 6) and qualitative (as shown in Figure. 5) results demonstrate their effectiveness.
The ablation study highlights the contributions of various components in our model, as summarized in Table 6. The full model (OUR’s) achieves the best performance with a Hausdorff Distance of 8.999 mm and a Chamfer Distance of 2.503 mm. Removing the mouth-convexity loss slightly degrades performance (0.11% in Hausdorff, 2.32% in Chamfer), while the asymmetry loss has a more noticeable impact (1.59% in Hausdorff, 2.72% in Chamfer). The latent code loss affects face consistency (0.53% in Hausdorff, 5.27% in Chamfer), and the geometry loss is critical, causing the largest increases among the four losses (2.64% in Hausdorff, 5.63% in Chamfer).
Data augmentation plays a pivotal role in model training, as its removal causes the most pronounced effect, increasing Hausdorff Distance by 5.75% and Chamfer Distance by 17.46%, while reducing the training dataset by 90%. The addition of augmented data appears to help reduce prediction error.
Figure 5 further underscores these findings, vividly illustrating the qualitative impact of each component. For instance, removing the mouth-convexity loss leads to an exaggerated chin, while the absence of the asymmetry loss results in pronounced facial asymmetry. Similarly, excluding the latent code loss causes notable drift in the cheek contours, and removing the geometry loss significantly exacerbates overall structural errors. Additionally, data augmentation seems to play an important role, as its absence introduces noticeable inconsistencies in facial predictions.
@@
Figure 5: Impact of losses and data augmentation on predictions. The left section illustrates actual surgical data, the middle section presents results after removing specific modules, and the right section showcases the full-model predictions.
4.2 Evaluation of Methodology Robustness
The results of our user study revealed that both engineers and medical professionals encountered similar challenges in accurately distinguishing between MLG faces and RSO. The minimal differences in sensitivity, specificity, and precision values between the two groups indicate that both groups faced similar difficulties, highlighting the reliability and accuracy of our novel method. The consistency in the performance of both groups confirms the effectiveness and robustness of our methodology and algorithm in predicting facial appearance following orthognathic treatment using only 3D face geometry.
4.3 Comparison with commercial VSP tools
Although Virtual Surgical Planning (VSP) tools (e.g., Dolphin, Morpheus 3D FaceMaker, and 3dMD Vultus) provide clinically accurate results for surgical planning, they require Computed Tomography (CT) or Cone Beam Computed Tomography (CBCT) imaging data as additional inputs [31].
A patient-oriented solution for visualizing surgical outcomes should be more accessible than a professional doctor-oriented one. In the early consultation stage, patients are not obligated to take CBCT or X-ray imaging, which would expose users to radiation. Instead, 3D facial reconstruction (e.g., 3dMD (3dMD Inc., Atlanta, GA, USA), Bellus3D (Bellus3D Inc., Los Gatos, CA, USA), and Ein-Scan3D (Shining 3D Technology, Hangzhou, China)) is much safer for post-treatment preview purpose.
@@ 4.3 Comparison with commercial VSP tools
Our fully automated and highly efficient method delivers accurate visualizations, offering an automatic preview of facial surgery outcomes that serves as an effective educational and motivational tool for patients, particularly during the consultation phase. By functioning as a visual aid, it enhances patients’ understanding and acceptance of treatment plans. To further engage patients, we generated an animation by interpolating the latent vectors, illustrating the transformation from pre-surgery to post-surgery. This visualization highlights the movement of the chin and the overall changes in facial appearance, as depicted in Fig. 6. For a more comprehensive demonstration, please refer to our supplementary video.
5 Limitation
Despite the comprehensive analysis conducted in this study, several limitations need to be acknowledged. The current training dataset is limited to Asian patients, which may impact the generalizability of the findings. Therefore, future research endeavors can focus on expanding the dataset to encompass greater diversity, including various ethnicities for improved applicability across different populations. Additionally, the study may benefit from addressing long-term skeletal changes, as the dataset primarily consists of adults with post-surgical scans taken at least six months after surgery to capture stabilized facial structures. While results are deemed reliable over an extended period, ongoing monitoring and consideration of potential long-term changes will be crucial for enhancing the study’s robustness. Furthermore, the model’s inability to offer multimodal-based predictions for individual patient attributes, such as age, gender, and skin condition, underscores the need for additional data with multi-modal labels to enable more targeted predictions without compromising data integrity. The current focus on providing automatic visualization services for patients rather than adjustable parameters for orthognathic professionals highlights a potential limitation for medical professionals seeking customizable features for surgical planning. In future work, considerations may include incorporating user-friendly interfaces with adjustable parameters to cater to professional needs effectively. Moreover, while the model primarily supports facial prediction for Asian facial types, efforts to expand its applicability to other ethnic groups through further research and validation are essential. Finally, while the subjective evaluation provided insights into the perceptual differences between MLG faces and RSO, it is crucial to recognise the limitations of relying solely on subjective assessments. Subjective measures, by nature, may introduce bias and subjective interpretation, potentially affecting the overall validity and reliability of the study findings. To mitigate these limitations, future studies can integrate quantitative clinical metrics and objective measures to complement subjective assessments, enhancing the study outcomes’ validity and reliability. Incorporating quantitative measures such as cephalometric analyses, facial landmark tracking, and patient-reported outcomes will facilitate a more comprehensive assessment and validation of machine learning-generated facial predictions in the context of orthognathic surgery.
6 Conclusions
In this paper, we introduce a novel method for predicting facial appearance following orthognathic treatment using only multi view images. During the training phase, our approach utilizes customized mouth-convexity and asymmetry losses, combined with latent code, geometric losses, and data augmentation, to enhance robustness and outperform existing methods in terms of accuracy.
Acknowledgments
This work was approved by the IRB (UW21-140 HKU/HA HKW IRB) and funded by the GRF grants (17107321) from the RGC of Hong Kong. The authors declare no competing interests. The trial is registered under HKUCTR-2971.